OpenStudy (anonymous):

solve log32 x=-2/5 solve logx 9=-1/2 solve log1/2 32=x

OpenStudy (cwrw238):

log32 x=-2/5 means that x = 32 ^ (-2/5) can you continue?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so x=.25?

OpenStudy (cwrw238):

yes 1 / (32^0.2)^2 = 1 / 2^2 = 1/4 0r 0.25

OpenStudy (anonymous):

ok thanks so much!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

so the second one 9=x^(-1/2) how do you get the answer?

OpenStudy (cwrw238):

9 = 1 --- x^(1/2) x^(1/2) = 1/9 now square both sides x = (1/9)^2

OpenStudy (anonymous):

do you do the same thing for the third one?

OpenStudy (cwrw238):

to get rid of the log function use : if loga b = x then b = a^ x or if you find this difficult to remember use : if log10 100 = 2 then 100 = 10^2 you now the last part is true so its a check that you have it right

OpenStudy (cwrw238):

for the third one you get (1/2)^x = 32 in this case take logs to the base e and use your calculator ln (0.5)^x = ln 32 x* ln(0.5) = ln 32 x = ln32 / ln(0.5)

OpenStudy (anonymous):

x=-5?